write a balanced nuclear equations for each of the following fusion reactions. write both the total mass of each side and the nuclear charge of each side.

Magnesium-24 is struck with a particle, producing chlorine-34 and a neutron.
Answer #1

Let us say that the particle = X

Given,

Total mass number on the left side of the equation = 24

Total mass number on the right side of the equation = (34 + 1) = 35

Hence, the mass number of X = (35 - 24) = 11

Total proton number on the left side of the equation = 12

Total proton number on the right side of the equation = (17 + 0) = 17

Hence, the proton number of X = (17 - 12) = 5

The element that contains 5 proton is boron with symbol B.

Hence, the symbol of X:

Therefore, the balanced equation of the fusion reaction: